# DiTAR: Diffusion Transformer Autoregressive Modeling for Speech Generation

### v2

@@ 3.4 Temperature for Continuous-Valued LMs
Table 1: Objective evaluation results of DiTAR and other systems on two subsets of LibriSpeech test-clean. Specifically, subset A is used and reported by NaturalSpeech3, while subset B is released by F5TTS. $\blacklozenge$ denotes the scores reported in NaturalSpeech3. $\clubsuit$ means the results obtained from the authors of F5TTS. $\spadesuit$ means the results are obtained via released checkpoints. The boldface and underline indicate the best and the second-best result, respectively. $\uparrow$ and $\downarrow$ indicate lower or higher values are better. Abbreviation: Disc.(discrete), Cont.(Continuous), AR(autoregressive model), NAR(non-autoregressive model), NFE(number of function evaluation)
@@
3.5 Implemenations
3.5.1 Continuous speech tokenization
Following LDM(Rombach et al., 2022), we use a variational auto-encoder (VAE)(Kingma, 2013) to convert the waveform into the distribution of latent $z$, represented by mean and variance. The encoder of the VAE consists of multiple layers of the convolutional network, and the decoder’s architecture follows BigVGAN(Lee et al., 2022). The adversarial training scheme also follows Lee et al. (2022). We adopt the multi-period discriminator (MPD) and multi-scale discriminator (MSD) proposed by Kong et al. (2020) as our discriminators. In our setup, the 24000hz waveform is compressed into 40hz latent with a dimension of 64.
3.5.2 Model
DiTAR consists of three modules: aggregation encoder, language model, and decoder (LocDiT). In our implementation, all of them are based on a transformer architecture. Specifically, both the encoder and decoder employ bidirectional attention masks, while the LM utilizes a causal attention mask. All transformers adopt the Pre-Norm(Xiong et al., 2020) architecture and utilize RMSNorm(Zhang & Sennrich, 2019) and RoPE (Su et al., 2024). Each patch of continuous tokens, together with a learnable special token positioned at the beginning of the sequence similar to (Devlin, 2018), is fed into the aggregation encoder. The output corresponding to the special token’s position serves as the aggregation embedding. Aggregation embeddings from different patches form a sequence that is processed by the LM. In LocDiT, LM’s outputs and the time embedding are added, along with historical context patches and noisy target tokens, forming a new sequence that serves as the input of LocDiT. When calculating the loss, we only consider the output corresponding to the position of noisy target tokens. During training, LM’s output is randomly replaced by a vector of all zeros with a probability of 0.1 to enable LM guidance for LocDiT.
@@ 3.5.2 Model
3.5.3 Diffusion formulation
Following Song et al. (2020b); Lu & Song (2024), we adopt a variance-preserving diffusion process defined by
$\bm{x}_{t}$ $=\alpha_{t}\bm{x}_{0}+\sigma_{t}\bm{\varepsilon}$ (5)
where $\bm{x}_{0}\thicksim q(\bm{x}_{0})$ denotes the data, $\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$ denotes the standard Gaussian noise, $t\in[0,1]$. We employ a conditional flow-matching loss(Lipman et al., 2022):
where the velocity is defined as: $\bm{v}(\bm{x}_{t},t)=\dot{\bm{x}_{t}}=\dot{\alpha}_{t}\bm{x}_{t}+\dot{\sigma}_{t}\bm{\varepsilon}$. At inference time, We employed the DDIM sampler(Song et al., 2020a), which is essentially an Euler ODE sampler with respect to signal-to-noise ratio $\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}}$ instead of $t$, proved to better align with the semi-linear property of the diffusion ODE(Lu et al., 2022a).
3.5.4 Zero-shot TTS system
The text sequence is converted into phonemes and processed through a lookup table to obtain text embeddings. Speech tokens are processed by the aggregation encoder to produce speech embeddings, which are then concatenated with the text embeddings. The embedding sequence serves as the input of the LM of DiTAR. During training, the loss of the text is not included. Additionally, we introduce a binary classifier with a fully connected layer at LM’s output to predict when to stop following Wang et al. (2017); Li et al. (2019). The loss function for zero-shot TTS can be summarized as $L=L_{diff}+L_{\text{stop}}$. During inference, prompting audio, text, and target text are fed as prefix input to DiTAR’s LM, which then autoregressively generates the target audio.
4 Experiments
4.1 SOTA Performance in Zero-Shot TTS
4.1.1 Setup
Many studies evaluate zero-shot TTS with inconsistent training sets, prompt audios, texts, and tools, diluting the value of comparisons. We standardize these variables for a more objective and fair assessment across systems.
Training and Evaluation Dataset. We consider two open-source datasets as our training dataset. 1) Librilight(Kahn et al., 2020), containing 60K hours of English speech data from LibriVox audiobooks. 2) Emilia(He et al., 2024), a multilingual dataset containing around 100k hours of speech data.
We adopt three open-source datasets for evaluation: 1) LibriSpeech(PC)(Panayotov et al., 2015; Meister et al., 2023) test-clean, containing 40 distinct English speakers and a 5.4-hour speech. We employ two established subsets: subset A from NaturalSpeech3, featuring 40 three-second speech prompts and 40 target samples, and subset B from F5TTS, which includes 40 prompts and 1127 samples. 2) Seed-ZH: a subset from DiDiSpeech 2(Guo et al., 2021), a Chinese speech dataset, containing 1088 prompts and targets. 3) Seed-EN: a subset from Common Voice(Ardila et al., 2019), a crowdsourcing English speech dataset with diverse accents, containing 2020 prompts and targets.
Evaluation Metrics. We evaluate four objective metrics: 1) Word Error Rate (WER), which assesses generation robustness. For consistency, we use the same ASR setups as previous studies to transcribe generated speech across different test sets. Specifically, we adopt a Hubert-based model [^1] https://huggingface.co/facebook/hubert-large-ls960-ft and Faster-whisper-large-v3  [^2] https://huggingface.co/Systran/faster-whisper-large-v3,version:0.10.1(Radford et al., 2022) for the subset A and B of Librispeech test-clean, respectively, Whisper-large-v3333https://huggingface.co/openai/whisper-large-v3 for Seed-EN and Paraformer-zh for Seed-ZH. 2) Speaker similarity to the prompt audio (SIM). Specifically, we employed WavLM-large(Chen et al., 2022) to compute the cosine distance between the generated and reference speech. 3) UTMOS(Saeki et al., 2022), an automatic predicted mean opinion score (MOS) to evaluate speech quality. 4) Floating point operations (FLOPs), measuring the computational load of models. The calculation process is detailed in A.5.
For subjective evaluation, we employ four MOS metrics: 1) N-MOS for naturalness, 2) Q-MOS for sound quality 3) S-MOS for speaker voice similarity 4) CMOS for side-by-side comparison with human audios.
Model Setup and Baselines. We benchmark DiTAR against diverse zero-shot TTS systems with varying architectures, including both multi-stage and single-stage schemes that operate in discrete or continuous value spaces. We conduct comparisons using DiTAR with 0.6 billion parameters and a patch size of 4. During inference, DiTAR’s LocDiT uses an NFE (Number of Function Evaluations) of 10. Specific details about the parameters of DiTAR are provided in A.2.
@@ 4.1.2 Experimental results
We conduct a multi-dimensional comparison of DiTAR with other baseline works. For objective metrics, Table 1 presents the evaluation results on LibriSpeech test-clean, while Table 6 in B.1 shows the results for Seed-ZH and Seed-EN. For subjective evaluation, we invite 10 English experts to rate the generated audio. For N-MOS, Q-MOS, and S-MOS metrics, the generated audios are rated on a scale of 1 to 5. For CMOS, experts compare the generated audio against the ground truth (GT) and assign scores from -2 to 2. Subjective results are detailed in Table 2.
Generation Robustness. As shown in Table 1, under two different training data configurations and test sets, DiTAR consistently delivered the best WER, showcasing robust synthesis performance matched by NAR systems with phone-level duration models. Table 6 in B.1 further details DiTAR’s comparison with models trained on proprietary data, highlighting its superior synthesis stability.
@@
Computaional Load. Table 1 indicates that DiTAR not only ensures high-quality audio generation but also dramatically cuts computational demands by approximately 343×3$\sim$ 43$\times$3 ∼ 43 × compared to other NAR systems. Although hybrid systems require less computational power, they tend to produce lower-quality outputs. Further details on inference efficiency are discussed in 4.4.2.
4.2 Scaling Behaviors
The language model and diffusion decoder benefit more from scaling. We scale the encoder, language model, and decoder (LocDiT) individually to assess their impacts, starting with a 0.4B parameter model. Given that the encoder and decoder process shorter sequences, we allocated more parameters to the language model empirically. Table 3 demonstrates that enlarging the language model and LocDiT enhances performance, whereas increasing the encoder size has little effect on outcomes.
Table 3: Scaling behavior of different components.
[…]
DiTAR (0.4B) 1.876 0.716
Encoder size ×$\times 4$ 1.821 0.72
Language model size ×$\times 4$ 1.695 0.727
LocDiT size ×$\times 4$ 1.785 0.726
Table 4: The impact of historical context in LocDiT. indictes that many samples fail to stop during generation.
[…]
1 2 3.334 0.716
1 6.131 0.692
0 (not used) 53 0.34
4 2 1.809 0.73
1 1.736 0.72
0 (not used) 22.874 0.56
Figure 3: The impact of patch size of LocDiT.
Figure 4: The impact of LM guidance under different NFE setups. $w=1$ indicates that guidance is not used.
Figure 6: The impact of temperature on generation diversity.
4.3 Method Analysis and Ablation Study
Patch Size. LocDiT utilizes bidirectional attention to generate the next patch. To investigate the impact of patch size, we vary it while keeping the model’s total parameter count constant. As illustrated in Figure 3, performance declines when patch sizes are either too large or too small. Excessively small patches diminish the model’s bidirectional attention capability, forcing reliance on causal-attention AR and degrading performance. This may explain the poor performance of causal AR with diffusion loss as noted in Li et al. (2024a). Conversely, overly large patches turn LocDiT into a bottleneck, necessitating increased parameters.
The Historical Context of LocDiT. We experiment with different numbers of historical patches and confirm the critical role of LocDiT’s context. Table 4 indicates that without historical context, the model’s performance drastically worsens. Integrating historical context shifts LocDiT’s function from mere generation to outpainting, markedly improving synthesis results. For larger patch sizes, such as 4, choosing a single historical context strikes the best balance between computational efficiency and performance.
LM Guidance. Figure 4 illustrates that LM guidance significantly enhances the diffusion decoder’s inference process. Without guidance ($w=1$), both WER and SIM deteriorate significantly. Conversely, an excessively large guidance scale can also impair outcomes. Additionally, even with extremely low NFE (such as 2), DiTAR still performs well on WER and SIM when combined with LM guidance.
4.4 Inference Analysis
4.4.1 The impact of temperature
To explore how $\tau$ balances diversity and stability, we synthesize the same text 500 times under different $\tau$ settings to assess voice diversity. We use only the text as DiTAR’s prefix input prompting the model to generate speech in random voices autoregressively. We then use WavLM-large to extract speaker embeddings from these samples. Following this, we conduct Principal Component Analysis (PCA), previously fitted on the training set, on the embeddings, and visualize the first two components. Figure 6 demonstrates that as $\tau$ increases, so does the diversity of the speaker voices. We also assess WER and SIM for zero-shot speech synthesis across different temperatures $\tau$, observing consistently excellent performance. Notably, higher temperatures slightly improve SIM scores, while lower temperatures enhance WER outcomes. Check more details in B.2.
4.4.2 Efficiency
DiTAR, blending a language model with a diffusion transformer, inherits features from both components. DiTAR drastically cuts computational demands by allocating the majority of its parameters to the language model and fewer to LocDiT.
We compare the inference performance of NAR (a diffusion transformer) and DiTAR under the same parameter sizes. During inference, all systems use CFG with NFE = 10. We evaluate throughput, defined as the duration of audio generated per unit time, across various batch sizes by inferring 10 seconds of audio on an A100 GPU. As shown in Fig 5, with small batch sizes, NAR achieves higher throughput due to lacking autoregressive computations. As batch sizes grow, NAR’s high FLOPs demands hinder throughput gains, while DiTAR’s throughput increases rapidly and is significantly superior to NAR.
5 Conclusion
In this work, we propose DiTAR which utilizes the diffusion transformer’s capacity for high-quality generation to create localized patches while maintaining the core autoregressive features of language models. For inference, we introduce temperature as the introduction time point for noise while solving the reverse diffusion ODE. Applied to zero-shot speech synthesis, DiTAR achieves SOTA robustness, speaker similarity, and naturalness with substantially lower computational requirements.
Appendix A Implementation Details
A.1 Training Details
Regarding data processing for LibriLight, since it lacks transcriptions, we use our internal ASR system to transcribe this dataset and then convert the text into phonemes for use. For Emilia, we perform Grapheme-to-Phoneme (G2P) conversion on the official transcriptions provided.
We utilize 16 A100 GPUs, each processing a batch size of 15K tokens, and train DiTAR for 0.5M steps. The AdamW optimizer is employed with a constant learning rate of 1e-4, $\beta_{t}=0.9$, and $\beta_{2}=0.99$. For DiTAR with 1B parameters, we utilize 32 A100 GPUs with a batch size of 7.5k per GPU.
A.2 Model Configuration for Scaling
During the validation of DiTAR’s scaling behavior, we trained models of four different sizes, ranging from 0.1 billion to 1 billion parameters. Specific hyperparameter configurations are detailed in Table 5.
Table 5: Configurations of DiTAR with different sizes.
Model size $\sim 0.1$B $\sim 0.4$B $\sim 0.6$B $\sim 1$B
Hyper-parameters Value
Encoder Number of layers 4 4 6 8
Hidden dim 512 1024 1024 1024
Number of heads 8 16 16 16
FFN dim 2048 4096 4096 4096
Language Model Number of layers 24 24 36 24
Hidden dim 512 1024 1024 1536
Number of heads 8 16 16 24
FFN dim 1024 4096 4096 6144
LocDiT Number of layers 4 4 6 8
Hidden dim 512 1024 1024 1024
Number of heads 8 16 16 16
FFN dim 2048 4096 4096 4096
A.3 Derivation of $x_{\theta}$ in temperature sampling for different parameterized diffusion
As previously discussed in Eq 5, $x_{\theta}$ denotes the predicted data, which can be derived under different parameterizations of diffusion. Consisdering a diffusion process defined by $x_{t}=\alpha_{t}x_{0}+\sigma_{t}\varepsilon$, where $x_{0}\thicksim q(x_{0})$ denotes the data, $\varepsilon\thicksim N(0,\textbf{I})$ denotes the standard Gaussian noise, $t\in[0,1]$.
For $\epsilon$-prediction mode,
$x_{\theta}(x_{t},t)=\frac{x_{t}-\sigma_{t}\epsilon_{\theta}(x_{t},t)}{\alpha_{t}}$ (8)
For $x_{0}$-prediction mode, $x_{\theta}(x_{t},t)$ is exactly model’s prediction.
For $v$-prediction mode, also known as flow-matching. Next, we will proceed with a step-by-step deduction. Based on the definition of $v$, we can derive the following:
$v$ $=\dot{\alpha_{t}}x_{0}+\dot{\sigma_{t}}\varepsilon$ (9)
$=\dot{\alpha_{t}}x_{0}+\frac{\dot{\sigma_{t}}(x_{t}-\alpha_{t}x_{0})}{\sigma_{t}}$ (10)
=($=\lparen\dot{\alpha_{t}}-\frac{\dot{\sigma_{t}}\alpha_{t}}{\sigma_{t}}\rparen x_{0}+\frac{\dot{\sigma_{t}}}{\sigma_{t}}x_{t}$ (11)
After rearanging $v$ and $x_{0}$, we get:
$x_{0}=\frac{\dot{\sigma_{t}}x_{t}-\sigma_{t}v}{\sigma_{t}\dot{\alpha_{t}}-\dot{\sigma_{t}}\alpha_{t}}$ (12)
Thus,
$x_{\theta}(x_{t},t)=\frac{\dot{\sigma_{t}}x_{t}-\sigma_{t}v_{\theta}(x_{t},t)}{\sigma_{t}\dot{\alpha_{t}}-\dot{\sigma_{t}}\alpha_{t}}$ (13)
A.4 The process of temperature sampling
Here, we present our temperature-based sampling process in the form of an algorithm flowchart, shown in Algorithm 1.
Algorithm 1 Temperature sampling
  $\eta\leftarrow\underset{n\in[1,N]}{\operatorname{argmax}}|t_{n}-\tau|$
  if $\eta=N$ then
     Sample initial noise $\bm{x}_{t_{N}}\sim\mathcal{N}(\bm{0},\bm{I})$
  else
     initial noise $\bm{x}_{t_{N}}\leftarrow\bm{0}$
  end if
  for $n=N-1$ to $1$ do
     $\hat{\bm{v}}\leftarrow\bm{f_{\theta}}(\bm{x}_{t_{N}},t_{N})$
     if $t_{n}=\eta$ then
        $\hat{\bm{x}}\leftarrow\mathcal{F}(\hat{\bm{v}})$
        Sample $\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})$
        $\bm{x}\leftarrow\alpha_{t_{n}}\hat{\bm{x}}+\beta_{t_{n}}\bm{z}$
     else
        $\bm{x}\leftarrow\bm{\Psi}(\bm{v},n+1,n)$
     end if
  end for
  Output: $\bm{x}$
A.5 Calculation of FLOPs
When calculating FLOPs for all systems, we use a 3-second audio prompt to synthesize 10 seconds of audio. Assuming a text frame rate of 7Hz, the prompt’s text length is 21, and the target text length is 70. We focus solely on the computationally dominant components, including causal transformers, non-causal transformers, and convolutional layers. For causal transformers, calculations account for the use of KV cache. The multi-head setup does not impact FLOPs, so we uniformly assume that the number of heads is 1. Considering that bias and normalization layers contribute minimally to the overall FLOPs of the transformer, we will temporarily disregard them. Matrix multiplication involves an equal number of additions and multiplications, so the result should be multiplied by 2. For the diffusion part, the corresponding FLOPs needs to be multiplied by the Number of Function Evaluations (NFE). Additionally, Classifier-Free Guidance (CFG) incurs double the computational cost.
For a one-dimensional convolution network with $N$ layers, a hidden channel size of $C$, a kernel size of $K$, and the input length of T, the FLOPs can be calculated as follows:
FLOPs $=[C,KT]\times[KT,C]\times N=2C^{2}KTN$ (14)
For a non-causal transformer with $N$ layers, a hidden size of $C$, an FFN intermediate hidden size of $C_{\text{mid}}$, and the input length of T, the FLOPs can be calculated as follows:
QKV $=[T,C]\times[C,3C]=[T,3C]=6TC^{2}$ (15)
Attention $=QK^{T}=[T,C]\times[C,T]=[T,T]=2T^{2}C$ (16)
V $=[T,T]\times[T,C]=[T,C]=2T^{2}C$ (17)
FC $=[T,C]\times[C,C]=[T,C]=2TC^{2}$ (18)
FFN_FC1 $=[T,C]\times[C,C_{\text{mid}}]=[T,C_{\text{mid}}]=2TCC_{\text{mid}}$ (19)
FFN_FC2 $=[T,C_{\text{mid}}]\times[C_{\text{mid}},C]=[T,C]=2TCC_{\text{mid}}$ (20)
[…]
For a causal transformer with $N$ layers, a hidden size of $C$, an FFN intermediate hidden size of $C_{\text{mid}}$, a prefix input length of $T_{\text{pre}}$, and the input length of T, the FLOPs can be calculated as follows:
[…]
QKV $=[1,C]\times[C,3C]=[1,3C]=6C^{2}$ (24)
Attention $=QK^{T}=[1,C]\times[C,t]=[1,t]=2t^{2}C$ (25)
V $=[1,t]\times[t,C]=[1,C]=2t^{2}C$ (26)
FC $=[1,C]\times[C,C]=[1,C]=2C^{2}$ (27)
FFN_FC1 $=[1,C]\times[C,C_{\text{mid}}]=[1,C_{\text{mid}}]=2CC_{\text{mid}}$ (28)
FFN_FC2 $=[1,C_{\text{mid}}]\times[C_{\text{mid}},C]=[1,C]=2CC_{\text{mid}}$ (29)
[…]
Appendix B Experimental Result Supplements
B.1 Comparison with additional models
We further compared DiTAR against a wider range of models, including some close-source ones, such as Seed-TTS. We conduct tests on datasets for two languages, namely Seed-EN and Seed-ZH. As shown in Fig 6, DiTAR achieves the best generation robustness and excellent speaker similarity. All results of other systems are reported in their papers.
@@ B.1 Comparison with additional models
B.2 The impact of temperature during inference

### v3

@@ 4.2 Scaling Behaviors
Table 1: Objective evaluation results of DiTAR and other systems on two subsets of LibriSpeech test-clean. Specifically, subset A is used and reported by NaturalSpeech3, while subset B is released by F5TTS. $\blacklozenge$ denotes the scores reported in NaturalSpeech3. $\clubsuit$ means the results obtained from the authors of F5TTS. $\spadesuit$ means the results are obtained via released checkpoints. The boldface and underline indicate the best and the second-best result, respectively. $\uparrow$ and $\downarrow$ indicate that lower or higher values are better. Abbreviation: Disc.(discrete), Cont.(Continuous), AR(autoregressive model), NAR(non-autoregressive model), NFE(number of function evaluation)
@@ 4.2.1 Experimental results
We conduct a multi-dimensional comparison of DiTAR with other baseline works. For objective metrics, Table 1 presents the evaluation results on LibriSpeech test-clean. For subjective evaluation, we invite 10 English experts to rate the generated audio. For N-MOS, Q-MOS, and S-MOS metrics, the generated audios are rated on a scale of 1 to 5. For CMOS, experts compare the generated audio against the ground truth (GT) and assign scores from -2 to 2. Subjective results are detailed in Table 2.
Generation Robustness. As shown in Table 1, under two different training data configurations and test sets, DiTAR consistently delivered the best WER, showcasing robust synthesis performance matched by NAR systems with phone-level duration models. Table 3 further details DiTAR’s comparison with models trained on proprietary data, highlighting its superior synthesis stability.
@@
Computaional Load. Table 1 indicates that DiTAR not only ensures high-quality audio generation but also dramatically cuts computational demands by approximately 343×3$\sim$ 43$\times$3 ∼ 43 × compared to other NAR systems. Although hybrid systems require less computational power, they tend to produce lower-quality outputs. Further details on inference efficiency are discussed in 4.5.2.
4.2.2 Comparison with Additional Models
To assess the upper-bound performance of DiTAR, we train DiTAR with 1 billion parameters on 280k-hour data and further compare DiTAR against a wider range of models, including some closed-source systems, such as Seed-TTS. We conduct tests on datasets for two languages, namely Seed-EN and Seed-ZH. As shown in Table 3, DiTAR achieves the best generation robustness and excellent speaker similarity. All results of other systems are reported in their papers.
@@ 4.2.2 Comparison with Additional Models
4.3 Scaling Behaviors
The performance of DiTAR consistently enhances as either data size or model size increases. We conduct scaling experiments concerning both data and model size. We expand the training data from 20k to 280k hours with a 0.6B model to assess performance changes. For model size, we scale from 0.1B to 1B parameters using 280k hours of training data, simultaneously increasing the parameters of the encoder, language model, and LocDiT. For the detailed setups, see Table 8. As shown in Figure 2, the model’s WER and SIM consistently improve as training data and model parameters increase.
The language model and diffusion decoder benefit more from scaling. We scale the encoder, language model, and decoder (LocDiT) individually to assess their impacts, starting with a 0.4B parameter model trained on 280k hours data . Given that the encoder and decoder process shorter sequences, we allocated more parameters to the language model empirically. Table 4 demonstrates that enlarging the language model and LocDiT enhances performance, whereas increasing the encoder size has little effect on outcomes.
Table 4: Scaling behavior of different components.
[…]
DiTAR (0.4B) 1.876 0.716
Encoder size ×$\times 4$ 1.821 0.72
Language model size ×$\times 4$ 1.695 0.727
LocDiT size ×$\times 4$ 1.785 0.726
4.4 Method Analysis and Ablation Study
Patch Size. LocDiT utilizes bidirectional attention to generate the next patch. To investigate the impact of patch size, we vary it while keeping the model’s total parameter count constant. As illustrated in Figure 3, performance declines when patch sizes are either too large or too small. Excessively small patches diminish the model’s bidirectional attention capability, forcing reliance on causal-attention AR and degrading performance. This may explain the poor performance of causal AR with diffusion loss as noted in Li et al. [34]. Conversely, overly large patches turn LocDiT into a bottleneck, necessitating increased parameters.
Figure 3: The impact of the patch size of LocDiT.
The Number of Historical Patches of LocDiT. We experiment with different numbers of historical patches and confirm the critical role of LocDiT’s context. Table 5 indicates that without historical context, the model’s performance drastically worsens. Integrating historical context shifts LocDiT’s function from mere generation to outpainting, markedly improving synthesis results. For larger patch sizes, such as 4, choosing a single historical context strikes the best balance between computational efficiency and performance.
Table 5: The impact of the number of historical patches in LocDiT. indicates that many samples fail to stop during generation.
[…]
1 2 3.334 0.716
1 6.131 0.692
0 (not used) 53 0.34
4 2 1.809 0.73
1 1.736 0.72
0 (not used) 22.874 0.56
LM Guidance. Figure 4 illustrates that LM guidance significantly enhances the diffusion decoder’s inference process. Without guidance ($w=0$), both WER and SIM deteriorate significantly. Conversely, an excessively large guidance scale can also impair outcomes. Additionally, even with extremely low NFE (such as 2), DiTAR still performs well on WER and SIM when combined with LM guidance.
Figure 4: The impact of LM guidance under different NFE setups. $w=0$ indicates that guidance is not used.
4.5 Inference Analysis
4.5.1 The impact of temperature
To explore how $\tau$ balances diversity and stability, we synthesize the same text 500 times under different $\tau$ settings to assess voice diversity. We use only the text as DiTAR’s prefix input prompting the model to generate speech in random voices autoregressively. We then use WavLM-large to extract speaker embeddings from these samples. Following this, we conduct Principal Component Analysis (PCA), previously fitted on the training set, on the embeddings, and visualize the first two components. Figure 5 demonstrates that as $\tau$ increases, so does the diversity of the speaker voices.